TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.MircoCoccoli on 2007-01-12 - 18:46

RAT Meeting in Point 8 - Jan 12th

PDF RAT_20070112.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x12x18x43/RAT_20070112.pdf

MircoCoccoli - 12 Jan 2007 - 18:45

Open Issues (Sect 78):

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 10.01.07 Connection of the NC Magnets in LSS8R (J-C.Guillaume)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-12 - 15:48

RAT for SCT in UA67/UJ67 - Jan, 12

PDF RAT_20070112.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x12x15x43/RAT_20070112.pdf

Open Issues:

  • 12.01 Power cut in ERD-R13
  • 12.01 Cleaning of filters by CV (planned for 17.01)
  • 10.01 QPS racks connection
  • 08.01 EE fan rotation problem to be fixed for the whole machine
  • 21.12 Aluminum panel for the cables ducts/ shipping and install.
  • 21.12 HV Insulation test of the DQS poles
Heat Runs and Test Schedule:
  • 08.01 600A and 13 kA EE system IST. ON GOING
  • 08.01 RB IST.
  • 11.01 120A powering tests.
  • 15.01 2 kA powering tests.
  • 17.01 PIC-PC-QPS Interface tests.
  • 19.01 600A powering tests.
  • 23.01 600A-EE powering tests.
  • 26.01 120A-600A-2kA 8h heat run.
  • 30.01 RQD&F powering tests.
  • 02.02 RB powering tests.
  • 07.02 RQ&RB 8h heat run.
  • 12.02 24h heat run test.
  • 14.02 AUG&UPS test.
  • 14.02 Polarity test.

-- Main.MirkoPojer - 12 Jan 2007 - 15:46

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-01-08 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login